OpenStudy (anonymous):

Help With 7th Grade Math

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

give me a question!

OpenStudy (anonymous):

Just give him a moment, he probably logged off

OpenStudy (anonymous):

Hi Im Back Im Sorry My computer Crashed

OpenStudy (anonymous):

\[\frac{ 3 }{ 5 }\times \frac{ 1 }{ 2}\]

OpenStudy (anonymous):

3/10

OpenStudy (anonymous):

10/21x3/5

OpenStudy (anonymous):

2/7